OpenStudy (anonymous):

If x ml pure alcohol is added to 400 ml of a 15% solution and after adding its strength becomes 32%, then the value of x is (A) 50 ml (B) 75 ml (C) 200 ml (D) 100 ml

OpenStudy (anonymous):

\[.15\times 400+x=.32(400+x)\]

OpenStudy (anonymous):

both sides representing the total amount of alcohol present

OpenStudy (anonymous):

Why .15×400+x=.32(400+x) ?

OpenStudy (anonymous):

I didn't understand the question.

OpenStudy (turingtest):

x%=x/100 the amount of alcohol in a given mixture is therefor given by the amount of liquid times the percent that is alcohol so the amount of alcohol in the original mixture is 15% of 400ml, or\[0.15\times400\]good so far?

OpenStudy (anonymous):

Yes, I understood

OpenStudy (turingtest):

ok, so if we add x ml of PURE alcohol, that is %100 concentration 100% X (ml of new mixture)=100% of x=x ml So we represent this as\[0.15\times400+1\times x=0.15(400)+x\]i.e. because it is pure alchohol we treat the coefficient of the amount of mixture being added as one. still good?

OpenStudy (anonymous):

Yes

OpenStudy (turingtest):

So then we are told our final mixture has 32% alcohol, and the number of milliliters of the substance now is 400+x ml so the amount of alcohol in our new mixture is then\[0.32(400+x)\]but this must be the same as the amount of alcohol we had before we mixed the two together, which was\[0.15(400)+x\] so they must be equal

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

100ml